OpenStudy (anonymous):

As x = (-π/2)-, sec(x) = ?

OpenStudy (danjs):

\[\sec(x) = \frac{ 1 }{ \cos(x) }\] \[\cos(-\pi/2)=\cos(\pi/2) = 0\]

OpenStudy (danjs):

secant of -pi/2 would be undefined.

OpenStudy (anonymous):

@DanJS So the point of reference is 90 degrees, with the point being (0,1). We are looking at the x value (cos)?

OpenStudy (danjs):

the angle here is \[\frac{ -\pi }{ 2 } ~~or~~ \frac{ 3\pi }{ 2 }\] The coordinate of the point on the unit circle corresponding to that angle is (x,y) = (0, -1)

OpenStudy (danjs):

Did you mistype some of the question? what is that - after the angle?

OpenStudy (anonymous):

It's as x = \[(\frac{ -\pi }{ 2 })-\] sex(x) = ?

OpenStudy (anonymous):

@DanJS